OpenStudy (lilsis76):

Please Help: How can I Prove the Following identities: a) tan^2 theta = sec^2 theta - 1 b) cot^2 theta = csc^2 theta -1

OpenStudy (lilsis76):

\[a) \tan^2\theta = \sec^2 -1\] \[b) \cot^2\theta = \csc^2\theta -1\] it says to use this identity as a guide: \[\sin^2\theta +\cos^2\theta = 1\]

OpenStudy (lilsis76):

@robtobey @nincompoop @aaronq can one of you please help me with these 2 problems. I dont understand what to do when i get to here: for a) ill work on the right side, and the same with b) a)\[\sec^2\theta - 1\] \[\rightarrow \frac{ 1 }{ \cos^2\theta } -1\] b) \[\csc^2\theta - 1\] \[\rightarrow \frac{ 1 }{ \sin^2\theta} -1\]

OpenStudy (lilsis76):

never mind I got it a friend just came over.